Compton Meeting, February 14, 2011

From PolWiki
Jump to: navigation, search

Agenda

Attendance

Kent, Dipangkar, Amrendra, Mark, Vladas, Wouter (minutes), Dave (chair)

Minutes

  • Only last few days we had good beam
    • before that, bad background, flucts in norm scint by factors
    • since Friday/Saturday we have better conditions, S/N of 50%
  • CPslow keeps locking up
    • memory issue? need to run memtest on it at next access
    • pictures of every lock-up would be useful
    • hcreboot switch from GEM will be used for Compton
    • Don set it up to reboot on exception error, not tested yet
  • Electron detector
    • presentation by Vladas (elog 151)
      • efficiency of ~8% for plane 1, ~2.5% for planes 2 and 3
      • efficiency-corrected population per strip looks reasonable
      • Compton-edge at strip 48, or 1.74 cm (according to Mark & Jay's calculation)
      • strip 0 is approximately half-way between rho = 0 and rho = 1
      • now 3.6 mm away from beam?? should be 8.5 mm to strip 0, so maybe 7 mm, but then guard strips, and first 4 strips are connected to ground... so maybe 6 mm
    • presentation by Amrendra (elog 150)
      • looks good
      • today will swap the cables
      • DD: different from Dec 10, now more hits on strips 0-32
      • should figure out what to do on 0-32, then repeat efficiency studies on this run
      • background-subtracted and background on hclog 220185, odd channels lower due to being read out on one side, should be taken into account when efficiency is calculated on these runs
    • Amrendra: 4mm away from beam, is this safe?
      • in IR, no large excursions when looking at EPICS during beam recovery
      • it has survived beam trips already, so probably OK
      • should have taken out for Moller measurement today probably
      • general feeling that it should be safe unless wild things happen
      • harp scans upstream? no difference with being in garage... so also acceptable
    • if access, need 10 mins to push QWADs
  • Laser
    • fine, apart for unhealthy CPslow
    • locking at 700W
    • time to go down? no, would need longer time
  • Photon detector
    • signal path, see elog 56
      • QDC is in data stream, only first QDC entry should correspond to the snapshot in the helicity event
      • threshold now 20mV for un-amplified signals coming up
    • shift of baseline
      • shifts in baseline from tunnel are large even in ms interval
      • pedestal shifts are not caused by the ADC itself, eliminated by Don's studies
      • phosphorescence in CsI(pure), see paper, evidence on 5-10s timescale
      • similar to what has been seen before at 20min timescale?
      • does correction as Don is doing right now taking care of this?
    • plans
      • Zener'ize the bases, characterize the tube, glue on to spare crystal
      • test: mini-megan, or detector group lab
      • build mini-megan into photon detector, Don has all the pieces
    • old tube still has the peak
      • probably not activation
    • need to evaluate background by looking at pedestal more carefully
    • amplifier in tunnel: what did we learn?
      • should set up amplifier system upstairs before moving it downstairs
  • Analysis
    • accum0 are not in MPS tree but are in HEL tree
    • QDC is not filled...